OpenStudy (anonymous):

Express in lowest terms X+X^4/X^2-X+1

OpenStudy (anonymous):

A; x(x + 1) B; 1 + x²

OpenStudy (anonymous):

C;X^4/2

OpenStudy (anonymous):

Is it the expression? \[\frac{ x+x^{4} }{ x^{2}-x+1 }\]

OpenStudy (anonymous):

YES

hartnn (hartnn):

hint : x^3+1 = (x+1)(x^2-x+1)

OpenStudy (anonymous):

\[\frac{ x+x^{4} }{ x^{2}-x+1 }=\frac{ x(1+x^{3}) }{ x^{2}-x +1}=\frac{ x(x+1)(x^{2}-x+1) }{ x^{2}-x+1 }=x(x+1)\]